A real estate developer offers to sell you some prime real estate for $584,000 today. You agree to pay $224,000 in exactly 6 months but the balance in exactly 22 months from today when you expect to receive some cash from an investment. How much will you need to pay the developer in 22 months if the interest rate is 13.5% per annum compounding monthly
Answer #1

PV of the real estate = 584,000

rate per month = 13.5%/12
PV of 224000 paid after 6 months = 224000/(1+13.5%/12)6 = 209,457.93

PV of amount paid after 22 months = 584000 - 209,457.93 = 374542.07
The amount after 22 months = 374542.07*(1+13.5%/12)22 = 479058.30

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well